OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (−2, −3) and has a slope of one half? 2x + y = 4 2x − y = 4 x + 2y = 4 x − 2y = 4

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@ankit042

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

using point-slope formula...\[(y-y_1)=m(x-x_1)\]where the values are\[m=\frac{1}{2}\]\[x_1=-2\]\[y_1=-3\]plug them in, and simplify in the form\[Ax+By+C=0\]this is the standard form of equation of a line....

OpenStudy (anonymous):

so -2+-3+1/3=

OpenStudy (anonymous):

no it's not... you plug-in the values in point-slope formula...

OpenStudy (anonymous):

\[(y-(-3))=\frac{1}{2}(x-(-2))\]

OpenStudy (anonymous):

all you have to do is simplify it to standard form...

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

take your time...

OpenStudy (anonymous):

2x − y = 4

OpenStudy (anonymous):

what's with ? in a diamond?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

2x − y = 4

OpenStudy (anonymous):

2 x − y = 4

OpenStudy (anonymous):

thats funny

OpenStudy (anonymous):

2x-y=4

OpenStudy (anonymous):

let us make it simple... (y+3)=(1/2)(x+2) 2(y+3)=x+2.... by cross multiplying 2 in left hand side 2y+6=x+2... by distributive property x-2y+2-6=0... transposing all in right hand side x-2y-4=0.... the answer in standard form

OpenStudy (anonymous):

im back

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

@Orion1213

OpenStudy (anonymous):

does this go with a b c or d

OpenStudy (anonymous):

it's D...

OpenStudy (anonymous):

thx man

OpenStudy (anonymous):

np...